# Modular Jump Gaussian Processes

### v1

@@ 5.3 Modified Michalewicz
For a higher-dimensional example, we adapted the so-called Michalewicz function, which can be found in the Virtual Library of Simulation Experiments (Surjanovic and Bingham,, 2013). Details can be found in Appendix A.3. Michalewicz is dimension-adaptive; here we present a 4d analysis to match the real data analysis coming in Section 5.4. An LHS design is used with $N=100,000$ training inputs. The marginal distribution of the response is shown in the center panel of Figure 6. Unlike the previous two data sets, the response is not plausibly composed of two Gaussian components. Instead, it is comprised of a point mass around 0.5 with other points scattered between in $[-4,0]$. Nevertheless, we find that EM with two Gaussian components works well in the Modular JGP setup. One goal of the presentation here is to demonstrate how even a naïve marginal clustering often works well in practice.
Figure 11: Left: MC experiment on the Modified Michalewicz data set; Right: A line plot showing the RMSE at individual runs for each of the four methods. The Modular JGP performs best in 27 of 30 runs.
Since $N$ here is very large, we must omit the global GP as a compactor in this experiment. Instead of a CGP we deploy a random forest as our classifier instead. Finally, we limit the number of MC reps to thirty. Results are provided in Figure 11. LAGP, optimal neighbors JGP, and the Modular JGP all considerably outperform the Original JGP. LAGP and OLAGP perform comparably, although interestingly LAGP seems to slightly outperform OLAGP in this case. The Modular JGP performs best of all, yielding the lowest RMSE in 27 of the 30 reps.
@@ 5.4 AMHV transport
Finally, consider our motivating AMHV transport data. Specifically, we used the larger version of the data set described in Section 5.2 of Park et al., (2023). The input dimension is $d=4$ and $N=9{,}500$. Although $N$ in moderately large in absolute terms, 4d is a large volume to fill. So, in a certain sense, this example is the most sparse of the ones entertained in this paper. Outputs are derived from a simulation of the behavior of AMHVs in a particular facility, tallying the average transport time of the AMHVs, including any time spent waiting for the vehicle to arrive. Park et al., (2023) show that as production in the facility increases, transport time, which depends on various design variables, increases and becomes a major bottleneck to the entire manufacturing process. The four varied here include vehicle speed, vehicle acceleration, minimum distance required between vehicles, and the maximum search range of empty vehicles.
A view of the marginal distribution of the response for these data was provided earlier, in the right panel of Figure 6. This is the first example we have looked at where the two clusters overlap. Additionally, the sparse nature of this data set means that local methods do not perform nearly as well. (The closest points in $X_{N}$ to a predictive location $x$ may actually be far and not representative.) This is a particular problem for OLAGP, which uses $x_{*}$ as a proxy for $x$. Consequently, we prefer a global rather than local GP in our application of the Modular JGP for these data.
Figure 12: Left: MC experiment on the AMHV transport data set; Right: A line plot showing the RMSE at individual runs for each of the five methods. The Modular JGP performs best in 95 of 100 runs.
The results of the experiment are shown in Figure 12. As expected, the local methods do not perform as well as the global ones, with OLAGP suffering most. But the global GP performs well, producing comparable results to the Original JGP. The Modular JGP produces only slightly better results, but is consistent in doing so (lowest RMSEs in 95 of the 100 reps).
@@ 6 Discussion
We introduced two ideas, adapting LAGP neighborhoods locally (OLAGP) learning a spatial feature from marginal clustering. Separately, both improve predictions for data from jump processes. They shine brightest when used together. We call this hybrid Modular JGP, since it involves many of the same ingredients as the Original JGP, except with off-the-shelf subroutines daisy-chained together.
One opportunity for future work pertains specifically to OLAGP. We mentioned previously that this method has thus far only been implemented for neighborhoods chosen using NN. However, it has been shown that NNs can be suboptimal compared to other subdesigns in the GP contexts (Vecchia,, 1988; Stein et al.,, 2004). Gramacy and Apley, (2015) demonstrated that neighborhoods chosen using via design criteria mean-square prediction error (MSPE) often yield superior LAGP predictions. We expect that to extend to the OLAGP context, but investigating that represents future work.
Other potential opportunities target the Modular JGP more holistically. The first involves extending the Modular JGP to noisy functions. The main concern in this case is whether the signal-to-noise ratio is large enough for the GP to accurately capture jumps in the data. The second involves extending the Modular JGP to allow more than two jump levels. Here we assumed that the marginal response of data from jump processes could be described as a mixture distribution with two components. However, it is possible for the response to have more than two levels. In that case, it might help to augment with a new feature for each added level. In other words, for $\ell$ levels we require $\ell-1$ new features. This would require a multiclass classification subroutine, however both of the methods we used – classification GP from kernlab and random forest from randomForest – have this capability.
@@ A.1 Phantom
The Phantom data set is generated in the same way as the Phantom data set in Park, (2022), but with different hyperparameter values. This version of the Phantom is evaluated on an evenly spaced $101\times 101$ grid in $[-0.5,0.5]^{2}$. The response is generated using the partitioning scheme shown in the left panel of Figure 13. In this plot, the color represents the true component membership of every input combination: white pixels belong to component 1, and black pixels belong to component 2. Data are drawn from a $\mathcal{N}_{N_{i}}(\mu_{i},\Sigma(X_{N_{i}}))$, where $N_{i}$ is the number of observations in component $i$, $\mu_{i}$ is the mean of component $i$, and $X_{N_{i}}$ is the subset of observations which belong to component $i$. The mean vector is $\mu=[27,0]^{\top}$ and the covariance matrix $\Sigma(X_{N_{i}})$ is calculated using the exponential covariance function (3) with $\tau^{2}=9$ and $\theta=[0.1,0.1]^{\top}$. After sampling all the data, the response was standardized so that $Y_{N}\in[0,1]$.
@@ A.2 Star
The Star data set is inspired by the Star data set in Park, (2022), but uses a different image for its partitioning structure. This version of the Star is evaluated on an evenly spaced $101\times 101$ grid in $[0,1]^{2}$. The response is generated using the partitioning scheme shown in the right panel of Figure 13. Like with the Phantom, the color represents the true component membership: white for component 1, and black for component 2. Data are drawn from a $\mathcal{N}_{N_{i}}(\mu_{i},\Sigma(X_{N_{i}}))$. The mean vector is $\mu=[-10,10]^{\top}$ and the covariance matrix $\Sigma(X_{N_{i}})$ is calculated using the exponential covariance function (3) with $\tau^{2}=3$ and $\theta=[0.01,0.01]^{\top}$. After sampling all the data, the response was standardized so that $Y_{N}\in[0,1]$.
@@ A.3 Modified Michalewicz
The Modified Michalewicz is a modification of the Michalewicz function from the Virtual Library of Simulation Experiments (Surjanovic and Bingham,, 2013). The original Michalewicz function is defined as $f(x)=-\sum_{i=1}^{d}\sin(x_{i})\sin^{2m}\bigl{(}\frac{ix_{i}^{2}}{\pi}\bigr{)}$ for $x\in[0,\pi]^{4}$ and $d$ in arbitrary dimension. We use $d=4$ and $m=10$ (the recommended default) with inputs coded into $[0,1]^{4}$. We first evaluate $f(x)$ using an LHS of size 100,000. The Modified Michalewicz is then defined as:
$y(x)=\begin{cases}f(x)&f(x)\leq-0.01\\ f(x)+0.5&f(x)>-0.01.\\ \end{cases}$ (7)
@@ Appendix B Global and LAGP Fits
We provide the mean fits given by a global GP and LAGP for the Phantom and Star data sets to highlight the improvements made by OLAGP and the Modular JGP. We also provide the difference between the predicted mean fit and the truth for each data set.
@@ B.1 Phantom
Modular JGP 0.1207
Figure 14: Left: Global GP fit. Center: LAGP fit. The blue dots are two predictive locations and the surrounding black dots are the selected neighborhood for that predictive location. Right: RMSE of Global GP, LAGP, OLAGP, and Modular JGP fits.
Both fits are visually less accurate than the OLAGP fit in Figure 4 and Modular JGP fit in Figure 7. This is verified by the table of RMSE values in the right panel, which shows that both OLAGP and the Modular JGP produce more accurate predictions than either method. The LAGP fit in the center panel also shows the selected neighborhoods for the same two predictive locations used in Sections 3 and 4, $(-0.32,0.17)$ and $(0.39,0.34)$. Notice in particular that the neighborhood for $(0.39,0.34)$ contains points in both the orange and yellow regions, which explains why our prediction is inaccurate. We are able to select smaller and differently shaped neighborhoods using OLAGP and Modular JGP so that our predictions are more accurate. We can verify that OLAGP and the Modular JGP make more accurate predictions by looking at Figure 15, which shows the absolute difference between the mean predictions and the truth.
Figure 15: Left: Difference between Global GP fit and truth; Center Left: Difference between LAGP fit and truth; Center Right: Difference between OLAGP fit and truth; Right: Difference between Modular JGP fit and truth.
@@ B.2 Star
Modular JGP 0.1462
Figure 16: Left: Global GP fit. Center: LAGP fit. The blue dots are two predictive locations and the surrounding black dots are the selected neighborhood for that predictive location. Right: RMSE of Global GP, LAGP, OLAGP, and Modular JGP fits.
Both fits are visually less accurate than the OLAGP and Modular JGP fits in Figure 9. The global GP performs particularly bad in this case, likely because the proportion of observations in component 1 is much larger than the proportion of values in component 2. The table of RMSE values in the right panel verifies our intuition, showing that OLAGP and the Modular JGP produce more accurate predictions than either method, particularly the global GP. The LAGP fit in the center panel also shows the selected neighborhoods for the same two predictive locations used in Section 5.2, $(0.15,0.8)$ and $(0.83,0.11)$. Notice in particular that the neighborhood for $(0.15,0.8)$ contains points in both the orange and yellow regions, which is why our prediction is inaccurate. We are able to select smaller and differently shaped neighborhoods using OLAGP and Modular JGP so that our predictions are more accurate. We can verify that OLAGP and the Modular JGP make more accurate predictions by looking at Figure 17, which shows the absolute difference between the mean predictions and the truth.
Figure 17: Left: Difference between Global GP fit and truth; Center Left: Difference between LAGP fit and truth; Center Right: Difference between OLAGP fit and truth; Right: Difference between Modular JGP fit and truth.

### v2

@@ 5.3 Modified Michalewicz
For a higher-dimensional example, we adapted the so-called Michalewicz function, which can be found in the Virtual Library of Simulation Experiments (Surjanovic and Bingham,, 2013). Details can be found in Appendix A.3; note that although the original function is in $0,\pi]^{4}$, we scale it to fall in $0,1]^{4}$. Michalewicz is dimension-adaptive; here we present a 4d analysis to match the real data analysis coming in Section 5.4. An LHS design is used with $N=100,000$ training inputs. The marginal distribution of the response is shown in the center panel of Figure 6. We cannot visualize the response surface in 4d, so instead we show a 2d version in the left panel of Figure 11.
Figure 11: Left: Modified Michalewicz response surface in 2d. Right: Image of AMHV transport device; taken from Park et al., (2023).
Unlike the previous two data sets, the response is not plausibly composed of two Gaussian components. Instead, it is comprised of a point mass around 0.5 with other points scattered between in [4,0][-4,0]. Nevertheless, we find that EM with two Gaussian components works well in the MJGP setup. One goal of the presentation here is to demonstrate how even a naïve marginal clustering often works well in practice.
Figure 12: Left: MC experiment on the Modified Michalewicz data set; Right: A line plot showing RMSPE at individual runs for each of the four methods. MJGP performs best in 29 of 30 runs.
Since $N$ here is very large, we must omit the global GP as a competitor in this experiment. Instead of a CGP we deploy a RF as our classifier. Finally, we limit the number of MC reps to thirty. Results are provided in Figure 12. LAGP, OLAGP, and MJGP all considerably outperform OJGP. LAGP and OLAGP perform comparably, although interestingly LAGP seems to slightly outperform OLAGP in this case. We suspect this is due to the spiky nature of the response surface; see Figure 11. MJGP performs best of all, yielding the lowest RMSPE in 29 of the 30 reps.
@@ 5.4 AMHV transport
Finally, consider the AMHV transport data first introduced in Section 1. Specifically, we used the larger version of the data set described in Section 5.2 of Park et al., (2023), scaled so that the inputs lie in the unit hypercube; $d=4$ and N=9,503N=9{,}503. A visual is provided in the right panel of Figure 11. Although $N$ is similar in magnitude to the sample size of the Star and Phantom data sets (N=10,201)(N=10{,}201), the addition of two dimensions means that there is a larger volume to fill. In that sense, this example is the most sparse of the ones entertained in this paper. Outputs are derived from a simulation of the behavior of AMHVs in a particular facility, tallying the average transport time of the AMHVs, including any time spent waiting for the vehicle to arrive. Park et al., (2023) show that as production in the facility increases, transport time, which depends on various design variables, increases and becomes a major bottleneck to the entire manufacturing process. The four variables here include vehicle speed, vehicle acceleration, minimum distance required between vehicles, and the maximum search range of empty vehicles.
A view of the marginal distribution of the response for these data was provided earlier, in the right panel of Figure 6. This is the first example we have looked at where the two clusters overlap. Additionally, the sparse nature of this data set means that local methods do not perform nearly as well. (The closest points in $X_{N}$ to a predictive location $x$ may actually be far and not representative.) This is a particular problem for OLAGP, which uses $x_{*}$ as a proxy for $x$. Consequently, we prefer a global rather than local GP in our application of MJGP for these data.
Figure 13: Left: MC experiment on the AMHV transport data set; Right: A line plot showing RMSPE at individual runs for each of the five methods. MJGP performs best in 99 of 100 runs.
The results of the experiment are shown in Figure 13. As expected, the local methods do not perform as well as the global ones, with OLAGP suffering most. But the global GP performs well, producing comparable results to OJGP. MJGP produces only slightly better results, but is consistent in doing so (lowest MSPEs in 99 of the 100 reps).
@@ 6 Discussion
We introduced two ideas, adapting LAGP neighborhoods locally (OLAGP) and learning a spatial feature from marginal clustering. Separately, both improve predictions for data from jump processes. They shine brightest when used together. We call this hybrid MJGP, since it involves many of the same ingredients as OJGP, except with off-the-shelf subroutines daisy-chained together.
One opportunity for future work pertains specifically to OLAGP. We mentioned previously that this method has thus far only been implemented for neighborhoods chosen using NN. However, it has been shown that NN neighborhoods can be suboptimal compared to other subdesigns in the GP contexts (Vecchia,, 1988; Stein et al.,, 2004). Gramacy and Apley, (2015) demonstrated that neighborhoods chosen via design criteria mean-square prediction error (MSPE) often yield superior LAGP predictions; see their paper for more information on neighborhoods chosen using MSPE. We expect their superior performance to extend to the OLAGP context, but investigating that represents future work.
Other potential opportunities target MJGP more holistically. The first involves extending MJGP to noisy functions. The main concern in this case is whether the signal-to-noise ratio is large enough for the GP to accurately capture jumps in the data. The second involves extending MJGP to allow more than two jump levels. Here we assumed that the marginal response of data from jump processes could be described as a mixture distribution with two components. However, it is possible for the response to have more than two levels. We believe MJGP could be easily extended to this case. The first two steps of clustering and classifying would be very similar to what we described in Section 4. First, cluster $Y_{N}$ to get $C_{N}=(c_{1},\dots,c_{N})^{\top}$ for $c_{i}\in\{1,\dots,\ell\}$, where $\ell$ is the total number of levels. Then use a multiclass classification method – both the CGP from kernlab and the RF from randomForest have this capability – to fit $\hat{f}(\cdot)=\mathrm{class}(X_{N},C_{N})$. Extract $\mathbb{P}(y(x)=j)$ for $j=1,\dots,\ell-1$. Let $\hat{f}_{j}(\cdot)=\mathbb{P}(y(x)=j)$ and use each to fit $\mathrm{GP}([\mathcal{X},\hat{f}_{1}(\mathcal{X}),\dots,\hat{f}_{\ell-1}(\mathcal{X})];[X_{N},\hat{f}_{1}(X_{N}),\dots,\hat{f}_{\ell-1}(X_{N})],Y_{N})$.
Although we limit our focus to GPs for computer experiments, there is extensive literature connecting GPs and geospatial statistics (Cressie,, 1993; Gelfand and Schliep,, 2016). Methods exist which fit GPs on errors after accounting for mean structure (Matheron,, 1969) or which can account for shifting local mean trends (Haas,, 1990). Methods even exist which similarly modify the distance-based kernels according to the covariance between points (Dumelle et al.,, 2023). Some of these approaches/problems are reminiscent of our jump processes. However, geostatistical methods are typically restricted to 2 or 3 inputs. Computer experiments often contain more input variables.
@@ A.1 Phantom
The Phantom data set is generated in the same way as the Phantom data set in Park, (2022), but with different hyperparameter values. This version of the Phantom is evaluated on an evenly spaced $101\times 101$ grid in $0.5,0.5]^{2}$. The response is generated using the partitioning scheme shown in the left panel of Figure 14. In this plot, the color represents the true component membership of every input combination: white pixels belong to component 1, and black pixels belong to component 2. Data are drawn from a $\mathcal{N}_{N_{i}}(\mu_{i},\Sigma(X_{N_{i}}))$, where $N_{i}$ is the number of observations in component $i$, $\mu_{i}$ is the mean of component $i$, and $X_{N_{i}}$ is the subset of observations which belong to component $i$. The mean vector is $\mu=[27,0]^{\top}$ and the covariance matrix $\Sigma(X_{N_{i}})$ is calculated using the exponential covariance function (3) with $\tau^{2}=9$ and $\theta=[0.1,0.1]^{\top}$. After sampling all the data, the response was standardized so that $Y_{N}\in[0,1]$.
@@ A.2 Star
The Star data set is inspired by the Star data set in Park, (2022), but uses a different image for its partitioning structure. This version of the Star is evaluated on an evenly spaced $101\times 101$ grid in $0,1]^{2}$. The response is generated using the partitioning scheme shown in the right panel of Figure 14. Like with the Phantom, the color represents the true component membership: white for component 1, and black for component 2. Data are drawn from a $\mathcal{N}_{N_{i}}(\mu_{i},\Sigma(X_{N_{i}}))$. The mean vector is $\mu=[-10,10]^{\top}$ and the covariance matrix $\Sigma(X_{N_{i}})$ is calculated using the exponential covariance function (3) with $\tau^{2}=3$ and $\theta=[0.01,0.01]^{\top}$. After sampling all the data, the response was standardized so that $Y_{N}\in[0,1]$.
@@ A.3 Modified Michalewicz
The Modified Michalewicz is a modification of the Michalewicz function from the Virtual Library of Simulation Experiments (Surjanovic and Bingham,, 2013). The original Michalewicz function is defined as $f(x)=-\sum_{i=1}^{d}\sin(x_{i})\sin^{2m}\bigl{(}\frac{ix_{i}^{2}}{\pi}\bigr{)}$ for $x\in[0,\pi]^{4}$ and $d$ in arbitrary dimension. We use $d=4$ and $m=10$ (the recommended default) with inputs coded into $0,1]^{4}$. We first evaluate $f(x)$ using an LHS of size 100,000. The Modified Michalewicz is then defined as:
[…]
@@ Appendix B Global and LAGP Fits
We provide the mean fits given by a global GP and LAGP for the Phantom and Star data sets to highlight the improvements made by OLAGP and MJGP. We also provide the difference between the predicted mean fit and the truth for each data set.
@@ B.1 Phantom
MJGP 0.1239
Figure 15: Left: Global GP fit. Center: LAGP fit. The blue dots are two predictive locations and the surrounding black dots are the selected neighborhood for that predictive location. Right: RMSPE of global GP, LAGP, OLAGP, and MJGP fits.
Both fits are visually less accurate than the OLAGP fit in Figure 4 and MJGP fit in Figure 7. This is verified by the table of RMSPE values in the right panel, which shows that both OLAGP and MJGP produce more accurate predictions than either method. The LAGP fit in the center panel also shows the selected neighborhoods for the same two predictive locations used in Sections 3 and 4, (0.32,0.17)(-0.32,0.17) and $(0.39,0.34)$. Notice in particular that the neighborhood for $(0.39,0.34)$ contains points in both the orange and yellow regions, which explains why our prediction is inaccurate. We are able to select smaller and differently shaped neighborhoods using OLAGP and MJGP so that our predictions are more accurate. We can verify that OLAGP and MJGP make more accurate predictions by looking at Figure 16, which shows the absolute difference between the mean predictions and the truth.
Figure 16: Left: Difference between global GP fit and truth; Center Left: Difference between LAGP fit and truth; Center Right: Difference between OLAGP fit and truth; Right: Difference between MJGP fit and truth.
@@ B.2 Star
MJGP 0.1512
Figure 17: Left: Global GP fit. Center: LAGP fit. The blue dots are two predictive locations and the surrounding black dots are the selected neighborhood for that predictive location. Right: RMSPE of global GP, LAGP, OLAGP, and MJGP fits.
Both fits are visually less accurate than the OLAGP and MJGP fits in Figure 9. The global GP performs particularly bad in this case, likely because the proportion of observations in component 1 is much larger than the proportion of values in component 2. The table of RMSPE values in the right panel verifies our intuition, showing that OLAGP and MJGP produce more accurate predictions than either method, particularly the global GP. The LAGP fit in the center panel also shows the selected neighborhoods for the same two predictive locations used in Section 5.2, $(0.15,0.8)$ and $(0.83,0.11)$. Notice in particular that the neighborhood for $(0.15,0.8)$ contains points in both the orange and yellow regions, which is why our prediction is inaccurate. We are able to select smaller and differently shaped neighborhoods using OLAGP and MJGP so that our predictions are more accurate. We can verify that OLAGP and MJGP make more accurate predictions by looking at Figure 18, which shows the absolute difference between the mean predictions and the truth.
Figure 18: Left: Difference between global GP fit and truth; Center Left: Difference between LAGP fit and truth; Center Right: Difference between OLAGP fit and truth; Right: Difference between MJGP fit and truth.
@@
Appendix C MJGP – Special Cases
Here we show the performance of MJGP on two special cases – data with small jumps and data with no jumps.
C.1 Data with Small Jumps
For the case with small jumps, we modify the function used in Section 1. We define the same jumping sine curve, but this time with a jump of 2 instead of 10. In other words, we use the function in Eq. 1, but replacing the 10 with a 2. We consider modeling this function using three of the four methods in Figure 1; we exclude MJGP with a global GP. The results are shown in Figure 19.
Figure 19: Predictive mean and 95% confidence intervals for the small jump case using a (left) global GP, (center) LAGP, and (right) MJGP.
The black lines represent the true function, while the colors (yellow, red, and blue) represent the predictive mean and 95% prediction intervals. We see that MJGP is still able to accurately capture the jumps, and offers improvement over the global GP and LAGP methods. Calculating RMSPE for each method confirms our findings; the global GP, LAGP, and MJGP produce RMSPEs of 0.8675, 0.4129, and 0.1038, respectively.
C.2 Data with No Jumps
For the case with no jumps, we use a sine curve defined on the same range as our previous functions (i.e., $x\in[0,100]$) to test MJGP. We use the same three methods for comparison as in the previous section. The results are shown in Figure 20.
Figure 20: Predictive mean and 95% confidence intervals for the no jump case using a (left) global GP, (center) LAGP, and (right) MJGP.
Here we see that the results do not look much different for any of the three methods. Since there are no jumps in the data, we do not expect using MJGP to offer any improvements. However, it is worth noting that using MJGP did not hurt us; the predictions made by MJGP are just as accurate as those made by LAGP. We confirm this statement using RMSPE; the global GP, LAGP, and MJGP produce RMSPEs of 0.0644, 0.0017, and 0.0015, respectively. This tells us that the method can be used when jumps are suspected but not confirmed without risk of overfitting.
Appendix D Computational Time
Table 2 shows the time (in minutes) of each of the simulations in Section 5.
Model Phantom Star Modified Michalewicz AMHV Transport
Global GP 4.92 3.75 N/A 12.87
LAGP 0.02 0.02 0.93 0.06
OLAGP 0.33 0.32 4.85 0.43
OJGP 22.32 22.22 352.95 30.57
MJGP 10.42 13.56 6.13 29.44
Table 2: Computation times for all data sets.
We report the computation time of a single Monte Carlo iteration since the computation time is not overly variable between reps. Times were recorded on an 8-core, Intel i7 CPU at 3.5GHz. We see that in general, the two JGP methods take the longest, with OJGP consistently taking the longest. This is especially apparent for the Modified Michalewicz function, although we recognize that this could likely be improved by parallelizing the code. We also see from this table that the time required for MJGP is highly dependent on the classification model used. Recall that part of MJGP is a GP fit, either global (AMHV Transport) or OLAGP (all other data sets). So we might expect MJGP to have similar computational efficiency to its corresponding GP method. However, this is not the case for all data sets. MJGP takes quite a bit longer to run than its GP counterparts for all data sets except the Modified Michalewicz function. This extra time is due to the classification model used; we see that MJGP takes longer, in general, when a CGP is used than when a RF is used.
Appendix E Prediction Interval Coverage
Figure 21 shows the prediction interval coverage for each of the simulations in Section 5.
Figure 21: Prediction interval coverage for the (left) Phantom, (left center) Star, (right center) Modified Michalewicz, and (right) AMHV transport data sets.
For each rep of each simulation, we recorded the percentage of true responses that fell within a 95% prediction interval of the predicted mean. We used those percentages to create the boxplots shown, with a dashed line at 95% to represent nominal coverage. For both the Phantom and Star data sets, shown in the left two panels, all methods except the global GP show coverage that is similar to each other but slightly higher than the nominal rate. Higher than nominal coverage indicates that prediction intervals are wider than necessary. This is a potential cause for concern, but since coverage is similar for all methods we prefer the method that produces the lowest RMSPEs, MJGP.
The results for the Modified Michalewicz and AMHV Transport data sets, shown in the right two panels, show a different trend. Here we see that all methods have lower than nominal coverage. Starting with the Modified Michalewicz data set, we see that OJGP, in addition to having the highest RMSPEs, also has the worst interval coverage of all methods. Similarly, OLAGP, which had the second highest RMSPEs, also has the second lowest interval coverage. LAGP, although having higher RMSPEs than MJGP, has better interval coverage. So LAGP may be preferable to MJGP for the Modified Michalewicz data set depending on the task at hand.
For the AMHV Transport data set, we again see that all methods undercover, but to a lesser extent than for the Modified Michalewicz data set. Here we see OLAGP produces the lowest interval coverage, which again matches the RMSPE trends we saw. We suspect this is caused by the aforementioned issues of OLAGP with sparse data. LAGP is less affected by these issues, showing better coverage than OJGP, which had the second lowest RMSPEs but second worst interval coverage. Both the global GP and MJGP have the best coverage, producing (on average) slightly lower the nominal coverage. Since MJGP also produced the lowest RMSPEs, we find that it was the best performing model overall.